Circular No. 5659

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

EU 1902+20 AND EU 2017–01 S. Brandt, A. J. Castro-Tirado, and N. Lund, Danish Space Research

Institute, communicate: “Two new x-ray transients, located at α = 19^{h}02^{m}, δ = +20^{o}.0 and α = 20^{h}17^{m}, δ = −1^{o}.0 (equinox 1950.0; estimated error radius 1^{o}), have been discovered with the WATCH wide-field monitor on the European Space Agency’s EURECA satellite. The new transients were detected in data from Nov. 22, but both appear at a lower level of significance already on Nov. 21. EU 2017–01 has a hard spectrum, whereas EU 1902+20 has only been detected between 6 and 8 keV, the flux from both sources being of the order 100 mCrab. Follow-up at other wavelengths is encouraged.”

PERIODIC COMET SWIFT-TUTTLE (1992t) J. Davies, Royal Observatory, Edinburgh, reports: “A 3.2- to 3.6-μm

spectrogram of P/Swift-Tuttle was taken at the U.K. Infrared Telescope on Nov. 13.15 UT by S. Lumsden on behalf of Davies, M. Mumma, S. Hoban, H. Weaver, and collaborators. Preliminary reduction of the spectrum shows a broad emission feature between 3.3 and 3.45 μm. There may be substructure within this feature, most notably appearing as a local peak near 3.427 μm. There is a broad emission feature centered at 3.52 μm, with no evidence of an emission feature at 3.28 μm. Detailed modeling of fluorescence from methanol produces a spectrum that explains all of the flux from 3.3 to 3.4 μm, and also that of the 3.52-μm feature. Excess flux is seen from 3.4 to 3.48 μm, probably due to another species. The 3.52-μm feature (which we attribute to methanol) has a flux of 8 × 10^{−16} W m^{−2} summed over an area 3′′.1 × 6′′.2, implying a methanol production rate of 1.5 × 10^{28} molecules s^{−1}. The flux in the 3.3- to 3.45-μm region is ∼ 6 × 10^{−15} W m^{−2} in excess of the underlying continuum. The underlying continuum has a flux of 1.15 × 10^{−13} W m^{−2} μm^{−1} near 3.25 μm, rising to 1.4 × 10^{−13} near 3.58 μm.”

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 Photometry by A. C. Gilmore with the 0.6-m f/14 Cassegrain reflector

at Mt. John University Observatory: Nov. 21.58 UT, V = 12.32 ± 0.06, U–B = –0.86 ± 0.07, B–V = +0.09 ± 0.06, V–R = +0.73 ± 0.07; Nov. 22.48, 12.39 ± 0.02, –0.74 ± 0.02, +0.05 ± 0.01, +0.74 ± 0.01.

1992 November 23 Daniel W. E. Green